OpenStudy (anonymous):

In a boarding kennel, dogs number 12 more than two times the number of cats. If there are 75 animals, how may cats are there?

OpenStudy (anonymous):

If cats are X, then dogs are 12 + 2 times the number of cats, or (2x + 12). 2x + 12 (dogs) + x (cats) = 75 3x + 12 = 75 3x = 63 x = 21 If 21 is the number of cats, 12 + 2*21 = 54

OpenStudy (anonymous):

Ooooh thanks, I have a hard time with word problems.

OpenStudy (anonymous):

No problem, all the best to you

OpenStudy (anonymous):

:)